# SpecGen: Automated Generation of Formal Program Specifications via Large Language Models

### v3

@@ I Introduction
  • A comprehensive evaluation from different aspects to evaluate the effectiveness of our approach. We compare SpecGen against Purely LLM-based approaches and representative non-LLM approaches. SpecGen succeeds in 279 out of the 385 programs, significantly outperforming the baseline approaches. Our code and data are available at [22].
@@ III-B2 Conversational Specification Generation
Given the initial prompt, LLM can obtain the initial output for the input program. As the output of the LLM in the first attempt may not successfully pass the validation of the verifier, we interleave the process of specification generation with verification failure feedback to prompt future generation in a conversational manner which is illustrated in Fig. 3. In particular, each generated specification by the model is verified by a JML verifier to test whether the generated result can pass the verifier. If the verification fails, we construct feedback information using the reported error message from the verifier as the prompt for the next generation. The verification error message can help the model understand the reason for failure and provide guidance for generating correct specifications. In addition, to avoid the verifier providing excessively long error messages, we configure the verifier to report only one verification failure message per attempt. Furthermore, through a massive amount of experiments, we summarize several types of common verification failures reported by the verifier. For each kind of error, we provide guidance in the natural language to facilitate the model in generating correct specifications. Detailed guidance information is provided on our website [22]. Upon encountering these types of verification failures reported by the verifier, we will insert corresponding guidance information into the prompt e.g., $<$Guidance$>$ in Fig. 3 to assist the model in resolving the issues. The conversation will be repeated iteratively until the specifications are successfully verified or a maximum number of iterations is reached.
@@ IV-A Implementation
We use the API provided by OpenAI [40] to communicate with the large language model of gpt-3.5-turbo-1106 for the experiments. We set the sampling temperature to 0.4 to balance the diversity and rigorousness of the outputs of GPT. 4 few-shot examples are used during the prompt construction to balance the input length and response time. The maximum number of rounds of conversation is set to 10. The verifier is OpenJML [26], one of the most recent JML specification verification tools, which checks the consistency between the given Java source code and the JML specifications within. Due to the incompleteness in the implementations of OpenJML, we set a timeout limit of 30 minutes for a single verification in our implementation to avoid unexpected situations, such as the non-responding of OpenJML. All experiments are conducted on an 8-core workstation with Intel Core i7-12700H CPU @2.30GHz and 32GB RAM, running Ubuntu 22.04.3 LTS. The version of OpenJDK is 1.8.0_371 for all experiments except for Houdini, which has to run under OpenJDK 1.6.0_45. We set the weight of comparative, logical, arithmetic, and predicative mutation to -1, -2, -4, and -4 respectively as the comparative mutation is more likely to pass the verification followed by the logical mutation. The predicative and arithmetic mutations are the least important through our observations from extensive experiments. Note that the weights are defined with negative values, leading to negative calculated scores as well. The reason for the design is to prioritize the specification candidates with fewer mutation occurrences.
@@ IV-B Dataset
To remedy this gap, we further collect another dataset, SpecGenBench, containing 120 samples as a supplement where 20 programs (including the corresponding specifications) from the dataset constructed by Nilizadeh et al. [43] and 100 programs from LeetCode [44]. This dataset is publicly accessible [22]. The selected programs are assured of the feasibility of expressing their behaviors as JML-specified verifiable specifications. These programs involve a variety of control flow structures and encompass multiple data structures such as arrays, strings, and other data structures supported by the Java library. They also cover a diverse set of specifications including post-conditions and loop invariants, involving both linear and nonlinear relationships between variables, making them representative of a broad spectrum of scenarios. They can be categorized into five categories according to their types of control flow structures [45]. Specifically, Sequential denotes the programs without branches or loops. Branched represents the loop-free programs that will contain branches like if-else or switch structures. Single-path Loop contains the simplest type of loop, with only one layer of loop structure without branches in their loop bodies. In contrast, Multi-path Loop denotes the loops that have branches in the loop bodies. Lastly, Nested Loop denotes the programs with multiple layers of loop structure where each layer may have a branch. The quantity of programs for Sequential, Branched, Single-path Loop, Multi-path Loop and Nested Loop is 26, 23, 24, 26, and 21, respectively. Programs in SpecGenBench have an average LoC of 20.77 and an average CC of 6.60.
@@ IV-C Baselines
AutoSpec [21]. It is a recent technique for specification generation combining LLMs and static analysis. The approach comprises three steps: code decomposition, hierarchical specification generation, and specification validation. AutoSpec first decomposes the input program into its components, upon which a hierarchy graph is built. For each component, AutoSpec queries the LLM for corresponding specifications respectively. Eventually, specifications for all components are combined to obtain the overall result, which is presented to a verifier for correctness verification. The progress is repeated iteratively until the result is successfully verified.
@@ IV-D Evaluation Metrics
User Rating. It aims to measure the semantic quality of the generated specifications. We invited 15 Ph.D. students who are experts in Java programming language to rate the specifications generated by different approaches. The research is conducted using the Likert Scale [49], where students are required to give a rating from one point to five points for each case. A reference rating criteria (can be accessed at [22]) is also presented to the students.
@@ V-A RQ1: Comparison with Baselines
Presented in Fig. 5 is the Venn diagram of the programs in SpecGenBench for which SpecGen and baseline approaches successfully generated verifiable specifications. It is noteworthy that SpecGen generates verifiable specifications for 7 programs that other baselines fail to yield, where 5 are from the Nested Loop, with the rest 2 from the Single-path Loop. Further investigation of these programs reveals that they are relatively complicated, making it hard for the baseline approaches to be correctly generated. The generated specifications of all approaches for these programs can be accessed on our website [22]. where a case study is also presented to illustrate the successful/unsuccessful generation.
@@ V-D RQ4: User Study on the Quality of Specifications
We can observe that the ground truth specifications (oracle) receive an average rating score of 4.83, indicating that the semantics of these programs can be described comprehensively through JML specifications. Furthermore, the specifications generated by SpecGen received a rating score of 4.54, which is close to the oracle, indicating that the generated specifications by SpecGen can also describe the real behaviors of the input program more fully. Among the 15 programs, all rating scores given to SpecGen are above 3, with the lowest rating being 3.33, meaning that in the worst case, SpecGen can still generate non-trivial specifications about the properties of the input program. In comparison, the specifications generated by Houdini and Daikon received an average rating score of 2.32, reflecting the semantic weakness in these specifications. Houdini and Daikon rely on pre-defined templates, which are in fact independent from the input program and can only cover a limited number of specification patterns. Consequently, the specifications produced are often simplistic and trivial, involving only a narrow range of variables and operators, making it difficult to capture the actual behavior and functionality of the input program precisely. Unlike traditional approaches that rely on a fixed set of templates, SpecGen utilizes the code comprehension capabilities of LLMs, which can cover a larger range of scenarios and generate targeted specifications that more closely match the semantics of the input program. We further provide a rating of 100 programs that SpecGen can successfully generate verifiable specifications following the criteria from Section IV-D on our website [22].

### v4

@@ I Introduction
  • A comprehensive evaluation from different aspects to evaluate the effectiveness of our approach. We compare SpecGen against Purely LLM-based approaches and representative non-LLM approaches. SpecGen succeeds in 279 out of the 385 programs, significantly outperforming the baseline approaches.
@@ III-B2 Conversational Specification Generation
Given the initial prompt, LLM can obtain the initial output for the input program. As the output of the LLM in the first attempt may not successfully pass the validation of the verifier, we interleave the process of specification generation with verification failure feedback to prompt future generation in a conversational manner which is illustrated in Fig. 3. In particular, each generated specification by the model is verified by a JML verifier to test whether the generated result can pass the verifier. If the verification fails, we construct feedback information using the reported error message from the verifier as the prompt for the next generation. The verification error message can help the model understand the reason for failure and provide guidance for generating correct specifications. In addition, to avoid the verifier providing excessively long error messages, we configure the verifier to report only one verification failure message per attempt. Furthermore, through a massive amount of experiments, we summarize several types of common verification failures reported by the verifier. For each kind of error, we provide guidance in the natural language to facilitate the model in generating correct specifications. Upon encountering these types of verification failures reported by the verifier, we will insert corresponding guidance information into the prompt e.g., $<$Guidance$>$ in Fig. 3 to assist the model in resolving the issues. The conversation will be repeated iteratively until the specifications are successfully verified or a maximum number of iterations is reached.
@@ IV-A Implementation
We use the API provided by OpenAI [39] to communicate with the large language model of gpt-3.5-turbo-1106 for the experiments. Temperature is set to 0.4 to balance the diversity and rigorousness of the outputs of GPT. 4 few-shot examples are used during the prompt construction to balance the input length and response time. The maximum number of rounds of conversation is set to 10. The verifier is OpenJML [25], the most recent JML specification verification tool to check the consistency between Java source code and JML specifications. Due to the incompleteness in the implementations of OpenJML, we set a timeout limit of 30 minutes for a single verification in our implementation to avoid unexpected situations, such as the non-responding of OpenJML. All experiments are conducted on an 8-core workstation with Intel Core i7-12700H CPU @2.30GHz and 32GB RAM, running Ubuntu 22.04.3 LTS. The version of OpenJDK is 1.8.0_371 for all experiments except for Houdini, which has to run under OpenJDK 1.6.0_45. We set the weight of comparative, logical, arithmetic, and predicative mutation to -1, -2, -4, and -4 respectively as the comparative mutation is more likely to pass the verification followed by the logical mutation. The predicative and arithmetic mutations are the least important through our observations from extensive experiments. Note that the weights are defined with negative values, leading to negative calculated scores as well. The reason for the design is to prioritize the specification candidates with fewer mutations.
@@ IV-B Dataset
To remedy this gap, we further collect another dataset, SpecGenBench, containing 120 samples as a supplement where 20 programs (including the corresponding specifications) from the dataset constructed by Nilizadeh et al. [42] and 100 programs from LeetCode [43]. The selected programs are assured of the feasibility of expressing their behaviors as JML-specified verifiable specifications. These programs involve a variety of control flow structures and encompass multiple data structures such as arrays, strings, and other data structures supported by the Java library. They also cover a diverse set of specifications including post-conditions and loop invariants, involving both linear and nonlinear relationships between variables, making them representative of a broad spectrum of scenarios. They can be categorized into five categories according to their types of control flow structures [44]. Specifically, Sequential denotes the programs without branches or loops. Branched represents the loop-free programs that will contain branches like if-else or switch structures. Single-path Loop contains the simplest type of loop, with only one layer of loop structure without branches in their loop bodies. In contrast, Multi-path Loop denotes the loops that have branches in the loop bodies. Lastly, Nested Loop denotes the programs with multiple layers of loop structure where each layer may have a branch. The quantity of programs for Sequential, Branched, Single-path Loop, Multi-path Loop and Nested Loop is 26, 23, 24, 26, and 21, respectively. Programs in SpecGenBench have an average LoC of 20.77 and an average CC of 6.60.
@@ IV-C Baselines
AutoSpec [21]. It is a recent technique for specification generation combining LLMs and static analysis. AutoSpec first decomposes the input program into its components, upon which a hierarchy graph is built. For each component, AutoSpec queries the LLM for corresponding specifications respectively. Eventually, specifications for all components are combined to obtain the overall result, which is presented to a verifier for correctness validation. The progress is repeated iteratively until the result is successfully verified.
@@ IV-D Evaluation Metrics
User Rating. It aims to measure the semantic quality of the generated specifications. We invited 15 Ph.D. students who are experts in Java programming language to rate the specifications generated by different approaches. The research is conducted using the Likert Scale [48], where students are required to give a rating from one point to five points for each case. A reference rating criteria is also presented to the students.
@@ V-A RQ1: Comparison with Baselines
Presented in Fig. 5 is the Venn diagram of the programs in SpecGenBench for which SpecGen and baseline approaches successfully generated verifiable specifications. It is noteworthy that SpecGen generates verifiable specifications for 7 programs that other baselines fail to yield, where 5 are from the Nested Loop, with the rest 2 from the Single-path Loop. Further investigation of these programs reveals that they are relatively complicated, making it challenging for the baseline approaches to be correctly generated.
@@ V-D RQ4: User Study on the Quality of Specifications
We can observe that the ground truth specifications (oracle) receive an average rating score of 4.83, indicating that the semantics of these programs can be described comprehensively through JML specifications. Furthermore, the specifications generated by SpecGen received a rating score of 4.54, which is close to the oracle, indicating that the generated specifications by SpecGen can also describe the real behaviors of the input program more fully. Among the 15 programs, all rating scores given to SpecGen are above 3, with the lowest rating being 3.33, meaning that in the worst case, SpecGen can still generate non-trivial specifications about the properties of the input program. In comparison, the specifications generated by Houdini and Daikon received an average rating score of 2.32, reflecting the semantic weakness in these specifications. Houdini and Daikon rely on pre-defined templates, which are in fact independent from the input program and can only cover a limited number of specification patterns. Consequently, the specifications produced are often simplistic and trivial, involving only a narrow range of variables and operators, making it difficult to capture the actual behavior and functionality of the input program precisely. Unlike traditional approaches that rely on a fixed set of templates, SpecGen utilizes the code comprehension capabilities of LLMs, which can cover a larger range of scenarios and generate targeted specifications that more closely match the semantics of the input program.
@@ VIII Conclusion
ACKNOWLEDGMENT
We are grateful for the constructive feedback of all the anonymous reviewers to improve this manuscript. The authors from Nanjing University are supported in part by the Leading-edge Technology Program of Jiangsu Natural Science Foundation (No. BK20202001), the National Natural Science Foundation of China (No. 62232008, 62172200), and the Postgraduate Research & Practice Innovation Program of Jiangsu Province (No. KYCX24_0237).